# Differentiable Fuzzy Cosmic-Web for Field Level Inference

### v1

@@ Appendix B Nonlocal Bias and Cosmic-Web Relation
Let us consider a scalar field $\eta$, which could be, for example, the matter overdensity field $\eta=\delta$ or the gravitational potential $\eta=\Phi$. We begin by constructing the Hessian matrix of this field, defined as:
@@ Appendix C Numerical Results of TEST2 and TEST3
Here we present the summary statistics resulting from TEST2 and TEST3. The case of TEST2–, when the resolution is doubled to 5 $h^{-1}\,\mathrm{Mpc}$ with respect to TEST1–yields similarly accurate results across all summary statistics (see 10, 11, and 12).
The additional numerical case TEST3 in which the bias parameters are assumed to be known, allows us to implement the most complex bias model as in Coloma-Nadal et al. (2024). Specifically, we define distinct bias prescriptions across the hierarchical cosmic-web, using a classification based on both the tidal field tensor and the Hessian of the density field. This results in 16 regions, each characterized by its own power-law, exponential threshold, and stochastic bias parameters. As shown in Figs. 13 14 and 15, the reconstruction maintains the same level of precision under this more sophisticated setting, as in TEST1 and TEST2.

### v3

@@ Appendix B Nonlocal bias and cosmic-web relation
Let us consider a scalar field $\eta$, which could be, for example, the matter overdensity field $\eta=\delta$ or the gravitational potential $\eta=\Phi$. We begin by constructing the Hessian matrix of this field, defined as
@@ Appendix C Numerical results of TEST2 and TEST3
Here we present the summary statistics resulting from TEST2 and TEST3. The case of TEST2 – when the resolution is doubled to 5 $h^{-1}\,\mathrm{Mpc}$ with respect to TEST1 – yields similarly accurate results across all summary statistics (see Figs. 10, 11, and 12).
The additional numerical case TEST3 in which the bias parameters are assumed to be known, allows us to implement the most complex bias model as in Coloma-Nadal et al. (2024). Specifically, we define distinct bias prescriptions across the hierarchical cosmic-web, using a classification based on both the tidal field tensor and the Hessian of the density field. This results in 16 regions, each characterised by its own power-law, exponential threshold, and stochastic bias parameters. As shown in Figs. 13 14, and 15, the reconstruction maintains the same level of precision under this more sophisticated setting, as in TEST1 and TEST2, assuming, however, the bias parameters to be known.